OpenStudy (anonymous):

What is the range of the graph of y = 3(x - 3)2 + 4?

OpenStudy (anonymous):

do you know the vertex for this parabola?

OpenStudy (anonymous):

no im taken a online class and i dont have much time to look it up and study because its due in 2 more days and i have a ton to do.

OpenStudy (anonymous):

do you at least know which way this parabola opens? upward or downward?

OpenStudy (anonymous):

if you just have to bust through with these problems, then use a computer algebra syste here is your example in wolfram. http://www.wolframalpha.com/input/?i=y+%3D+3%28x+-+3%29^2+%2B+4 you can see that it says the minimum is 4, so the range is \([4\infty)\) which you can read right off of the equation since \[(x-3)^2\geq 0\] for all x, since it is a perfect square, and so \[(x-3)^2+4\geq 4\]